OpenStudy (anonymous):

plz help me how to do i simpligy this -2 (index 5 ) √160s^2t^6

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

plz i have a huge test tomorrow help

ganeshie8 (ganeshie8):

\(\huge -2\sqrt[5]{160s^2t^6}\)

ganeshie8 (ganeshie8):

its like that ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it says simpilfy it how plz help

ganeshie8 (ganeshie8):

yeah we can simplify it,

ganeshie8 (ganeshie8):

we can pull out some t, and 160

OpenStudy (anonymous):

can you show me step by step how to do it

ganeshie8 (ganeshie8):

lets work on \(t\) first \(\huge -2\sqrt[5]{160s^2t^6} \) \(\huge -2\sqrt[5]{160s^2t^{5+1}} \)

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

u knw this exponent property : \(a^{m+n} = a^m.a^n\) ?

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

lets work on \(t\) first \(\huge -2\sqrt[5]{160s^2t^6} \) \(\huge -2\sqrt[5]{160s^2t^{5+1}} \) \(\huge -2\sqrt[5]{160s^2t^5t^1} \)

ganeshie8 (ganeshie8):

fine ?

ganeshie8 (ganeshie8):

nw we can pull out t^5, it becomes t

ganeshie8 (ganeshie8):

lets work on \(t\) first \(\huge -2\sqrt[5]{160s^2t^6} \) \(\huge -2\sqrt[5]{160s^2t^{5+1}} \) \(\huge -2t\sqrt[5]{160s^2t^1} \)

OpenStudy (anonymous):

why did u pull out t5 why not the whole t6

ganeshie8 (ganeshie8):

lets work on \(t\) first \(\huge -2\sqrt[5]{160s^2t^6} \) \(\huge -2\sqrt[5]{160s^2t^{5+1}} \) \(\huge -2\sqrt[5]{160s^2t^5t^1} \) \(\huge -2t\sqrt[5]{160s^2t^5t} \)

ganeshie8 (ganeshie8):

good question :)

OpenStudy (anonymous):

:)

ganeshie8 (ganeshie8):

we can pull out only in powers of 5, coz the t is under 5th index ha ?

OpenStudy (anonymous):

ahh ok thanks go on

ganeshie8 (ganeshie8):

so t^5 , when it comes out, becomes t

OpenStudy (anonymous):

ok i uderstand

ganeshie8 (ganeshie8):

we are half done, lets work on 160

ganeshie8 (ganeshie8):

160 = 16 x 10

ganeshie8 (ganeshie8):

160 = 16 x 10 160 = 16 x 2 x 5

ganeshie8 (ganeshie8):

160 = 16 x 10 160 = 16 x 2 x 5 160 = 4 x 4 x 2 x 5

ganeshie8 (ganeshie8):

we trying to see if we can make the exponent 5 , so that we pull out of radical something

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

160 = 16 x 10 160 = 16 x 2 x 5 160 = 4 x 4 x 2 x 5 \(160 = 2^2 \times 2^2 \times 2 \times 5\)

ganeshie8 (ganeshie8):

see any exponent 5 ?

OpenStudy (anonymous):

what is exponent sorry i forgot

ganeshie8 (ganeshie8):

\(a^m\) : \(m\) is the exponent

ganeshie8 (ganeshie8):

exponent is also called power

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

its the thing on top

OpenStudy (anonymous):

i understand

OpenStudy (anonymous):

then what

ganeshie8 (ganeshie8):

160 = 16 x 10 160 = 16 x 2 x 5 160 = 4 x 4 x 2 x 5 \(160 = 2^2 \times 2^2 \times 2 \times 5\)

ganeshie8 (ganeshie8):

do u see , any power = 5

OpenStudy (anonymous):

ye 2 2 and 1

ganeshie8 (ganeshie8):

160 = 16 x 10 160 = 16 x 2 x 5 160 = 4 x 4 x 2 x 5 \(160 = 2^{2+2+1} \times 5\)

ganeshie8 (ganeshie8):

Excelelnt !

ganeshie8 (ganeshie8):

160 = 16 x 10 160 = 16 x 2 x 5 160 = 4 x 4 x 2 x 5 \(160 = 2^{2+2+1} \times 5\) \(160 = 2^5\times5\)

ganeshie8 (ganeshie8):

so, lets put this in our original expression

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

\(\huge -2t\sqrt[5]{160s^2t^5t} \)

ganeshie8 (ganeshie8):

\(\huge -2t\sqrt[5]{2^55 s^2t^5t} \)

ganeshie8 (ganeshie8):

we can pull 2^5 out of radical, it becomes ?

OpenStudy (anonymous):

32?

ganeshie8 (ganeshie8):

no, 2^5 out of radical, it becomes just 2

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

\(\huge -2t \ 2\sqrt[5]{5 s^2t^5t} \)

OpenStudy (anonymous):

wait are u on facebook

ganeshie8 (ganeshie8):

\(\huge -4t\sqrt[5]{5 s^2t} \)

ganeshie8 (ganeshie8):

thats our final simplified expression,

OpenStudy (anonymous):

ok thanx

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

u dont like facebook

ganeshie8 (ganeshie8):

there, t^5 is redundantly coming inside radical, even after we simplified t, plz ignore that :)

ganeshie8 (ganeshie8):

oh gimme ur id, il add :)

OpenStudy (anonymous):

umarnaveed1@hotmail.com

OpenStudy (anonymous):

how old are u